4x=8y
2x+5y=27 solve by substitution

Answer by rfer(16322)

4x=8y
x=8y/4
x=2y
--------------
2x+5y=27
2(2y)+5y=27
9y=27
y=27/9
y=3
--------------
2x+5(3)=27
2x=27-15
x=12/2
x=6
---------------
(6,3)
--------------
x=2y
6=2*3
6=6